# DeepCell: Multiview Representation Learning for Post-Mapping Netlists

### v1

@@ Abstract
Representation learning for post-mapping (PM) netlists is a critical challenge in Electronic Design Automation (EDA), driven by the diverse and complex nature of modern circuit designs. Existing approaches focus on intermediate representations like And-Inverter Graphs (AIGs), limiting their applicability to post-synthesis stages. We introduce DeepCell, a multiview representation learning framework that integrates structural and functional insights from both PM netlists and AIGs to learn rich, generalizable embeddings. At its core, DeepCell employs the novel Mask Circuit Modeling (MCM) mechanism, which refines PM netlist representations in a self-supervised manner using pretrained AIG encoders. DeepCell sets a new benchmark in PM netlist representation, outperforming existing methods in predictive accuracy and reconstruction fidelity. To validate its efficacy, we apply DeepCell to functional Engineering Change Orders (ECO), achieving significant reductions in patch generation costs and runtime while improving patch quality.
Index Terms:
circuit representation learning, mask modeling, multiview learning, engineering change order (ECO)
§§footnotetext: Both authors contributed equally to this research.
@@ I Introduction
Representation learning has emerged as a powerful paradigm across domains such as computer vision (CV) and natural language processing (NLP), where pretrained models finetuned for specific tasks achieve state-of-the-art results. Inspired by these advancements, the field of Electronic Design Automation (EDA) has explored similar methodologies for circuit representation learning. Existing approaches, such as DeepGate Family [1, 2], FGNN [3], Gamora [4] and HOGA [5], demonstrate significant improvements in tasks like testability analysis, circuit identification, and design verification by focusing on And-Inverter Graphs (AIGs). However, this narrow reliance on AIGs limits their generalization to more complex and practical circuit abstractions.
Post-mapping (PM) netlists, composed of diverse standard cells, represent a key stage in the design flow but remain underexplored in circuit representation learning. These netlists introduce challenges due to their structural and functional heterogeneity, which is difficult to be effectively captured by existing AIG-based methods. Current solutions, such as simulation-based supervised training [2] or contrastive learning [3], struggle with scalability and efficiency in handling PM netlists. This gap hinders progress in critical post-mapping tasks like technology mapping and functional Engineering Change Orders (ECO).
@@
To address these challenges, we propose DeepCell, a novel multiview representation learning framework for PM netlists. DeepCell integrates information from both PM netlists and AIGs using a Graph Neural Network (GNN)-based PM encoder and a pretrained AIG encoder. At its core, DeepCell employs Mask Circuit Modeling (MCM), a self-supervised mechanism inspired by Masked Language Modeling (MLM) [6], which leverages AIG embeddings to refine PM netlist representations. By bridging the structural-functional gap inherent in PM netlists, DeepCell achieves rich and generalizable embeddings. We then validate DeepCell through its application to the functional ECO task, a critical post-mapping challenge involving design modification after tape-out [7]. Integrated as a plug-in to an existing ECO tool, DeepCell significantly reduces patch generation costs and runtime.
Our contributions are summarized as follows:
  • We propose DeepCell, the first multiview and multimodal representation learning framework tailored for PM netlists, integrating structural and functional insights from diverse standard cells.
  • We introduce Mask Circuit Modeling (MCM), a self-supervised mechanism for refining PM netlist embeddings using AIG-based representations.
  • We demonstrate the utility of DeepCell in functional ECO, achieving reductions in patch generation costs, gate count, and runtime while maintaining high-quality results.
@@
The remainder of this paper is organized as follows: Section II reviews related work. Section III describes the proposed DeepCell framework, including architecture and Mask Circuit Modeling mechanism. Section IV presents the pretraining results and investigate the effect of proposed training strategy. Next, we apply DeepCell in functional ECO tasks. Finally Section VI concludes this paper.
Figure 1: Overview of DeepCell
@@ II-A Circuit Representation Learning
Circuit representation learning has emerged as an attractive direction in the field of EDA, focusing on training models to obtain general circuit embeddings that can be applied to various downstream tasks [8]. The first circuit representation learning framework, DeepGate [1], proposes supervising model training using logic-1 probability under random simulation and achieves substantial improvements in tasks like testability analysis [9] and power estimation [10]. Its successor, DeepGate2 [2], further refines this approach by separating functional and structural embeddings for different applications. Additionally, Gamora [4] and HOGA [5] leverage sub-circuit identification as pre-training tasks, while FGNN [3] trains models in an unsupervised contrastive manner by distinguishing between equivalent and non-equivalent circuits. Despite these advancements, existing models primarily focus on learning representations of AIG netlists. There is still no available solution capable of learning post-mapping netlists with arbitrary logic cells and addressing the practical applications on post-mapping stages.
While these advancements primarily address representation learning for intermediate formats such as AIGs, the challenges posed by post-mapping netlists remain largely unexplored. Functional ECO represents one such critical post-mapping application, where effective representation learning could significantly enhance performance and efficiency.
@@ II-B Functional ECO
ECO are a critical component in the VLSI design process, used to rectify design problems after tape-out or chip fabrication. ECO involve making modifications to correct these errors, and they are indispensable in avoiding the high expenses associated with design re-spin  [11]. For functional ECO, the purpose is to generate patch so that the original circuit is equal to the golden circuit, while minimizing the resource cost of the generated patches and making the running time as short as possible. Synthesis-based ECO algorithms are good at solving this problem [7]. It relies on a diagnostic strategy to identify internal rectifier signals, and then applies a resynthesis technique to generate patch functions for functional differences. These algorithms have been able to automate the process of functional ECO.
@@ III-A Overview
Fig. 1 presents the overview of DeepCell framework, which consists of an AIG encoder, a PM encoder and a Transformer model. The framework operates in two stages to capture the general representation of post-mapping (PM) netlists. In Stage 1, the PM encoder is trained to learn cell embeddings (Cell Emb.) from the PM netlist. In Stage 2, the AIG encoder, which has been pretained and frozen, is utilized to extract gate embeddings (Gate Emb.) from the corresponding AIG netlist. Then, we mask a random subset of the cell embeddings and reconstruct these masked embeddings using the gate embeddings.
@@ III-B1 Node Features
PM netlists consist of a wide variety of standard cells, making it impractical to represent them using one-hot encoding for each type of cell. Therefore, To address this, we embed the truth table of each standard cell into its corresponding node feature $x_{i}$. Formally, the node feature encoding is defined in in Eq. (1), where $D$ is the dimension of node feature vector and $tt_{i}$ represents the 0/1 truth table vector of standard cell $i$. The truth table $tt_{i}$ is repeated until the node feature vector reaches the specified dimension $D$. In the default setting, we assign $D=64$, ensuring that this encoding mechanism is adaptable to PM netlists across various technology libraries and supports arbitrary logic units with up to 6 inputs.
@@
For example, the cell xor2_1 defines XOR functionality with 2 inputs and 1 output. Its truth table, 0110, is extracted from the technology library and expanded into a 64-dimensional feature vector by repeating the pattern. Thus, the node feature of the xor2_1 cell becomes 0110 0110 ... 0110.
@@ III-B2 Aggregator
For structural embedding aggregation, $aggr^{s}$ is implemented using the GCN aggregator [12], which aggregates messages from the predecessors of $i$. Here, $\mathcal{P}(i)$ denotes the set of fanin cells of $i$:
@@
For functional embedding aggregation, $aggr^{f}$ is implemented using a self-attention aggregator [13] to distinguish the functionality of the predecessors. Unlike AIG netlists, which consist solely of AND gates and inverters, PM netlists contain diverse standard cells. To account for this diversity, we differentiate cells using their node features $x_{i}$ and introduce an update function, $update$. Formally, the functional aggregation process in DeepCell is defined as Eq. (3), with the $update$ function implemented as a multi-layer perceptron (MLP).
@@
Finally, the embeddings of PM netlist are denoted as Eq. (4), where $encoder$ is the above GNN-based PM encoder.
$\begin{split}\mathbf{H}^{P}&=encoder(\mathcal{G}^{P})\\ &H_{i}^{P}\in\mathbf{H}^{P},i\in\mathcal{V}^{P}\end{split}$ (4)
@@ III-C AIG Encoder
Our DeepCell framework incorporates a multiview representation learning mechanism, enabling it to learn cell embeddings in PM netlists from an additional perspective provided by AIGs. Specifically, given a PM netlist $\mathcal{G}^{P}=(\mathcal{V}^{P},\mathcal{E}^{P})$, we convert it into an AIG netlist $\mathcal{G}^{A}=(\mathcal{V}^{A},\mathcal{E}^{A})$. We then employ DeepGate2 [2] as the AIG encoder to derive the gate-level embeddings $H_{j}^{A}\in\mathbf{H}^{A}$ as formulated in Eq. (5).
$\begin{split}\mathbf{H}^{A}&=DeepGate2(\mathcal{G}^{A})\\ &H_{j}^{A}\in\mathbf{H}^{A},j\in\mathcal{V}^{A}\end{split}$ (5)
@@ III-D Mask Circuit Modeling
Figure 2: Mask Circuit Modeling
We refine the cell embeddings $\mathbf{H}^{P}$ using information from the AIG view by mask circuit modeling (MCM). As shown in Fig. 2, given a PM netlist $\mathcal{G}^{P}$, we convert $\mathcal{G}^{P}$ to AIG Netlist $\mathcal{G}^{A}$. Both $\mathcal{G}^{P}$ and $\mathcal{G}^{A}$ are encoded by PM encoder and AIG encoder in Eq. (4) and Eq. (5), respectively.
To apply MCM, we randomly select node $p\in\mathcal{G}^{P}$ and extract a $k$-hop predecessors around node $p$, denoted as the masked hop $\mathcal{M}(p)$. Within the masked hop $\mathcal{M}(p)$, we replace the functional embeddings $hf_{i}$ of the nodes with a learnable masked token $hm$, while preserving their structural embeddings $hs_{i}$.
Therefore, we totally have $M+N$ tokens (see Eq. (6)), where $M=|\mathcal{V}^{P}|$ and $N=|\mathcal{V}^{A}|$. All the selected masked nodes $p$ is in set $\mathcal{S}$.
$\begin{split}{H_{i}^{P}}&=\{hs_{i}^{P},hf_{i}^{P}\},~{}i\in\mathcal{V}^{P},i\notin\mathcal{M}(p)\\ {H_{i}^{P}}&=\{hs_{i}^{P},hm\},~{}i\in\mathcal{V}^{P},i\in\mathcal{M}(p)\\ {H_{j}^{A}}&=\{hs_{j}^{A},hf_{j}^{A}\},~{}j\in\mathcal{V}^{A}\\ \end{split}$ (6)
We process these $M+N$ tokens using a Transformer model $\mathcal{T}$. Formally, we define the input and output of the Transformer model as below. The training objective is to reconstruct the masked area using the remaining cell embeddings and the learned gate embeddings from AIG view.

### v2

# DeepCell: Self-Supervised Multiview Fusion for Circuit Representation Learning

@@ Abstract
We introduce DeepCell, a novel circuit representation learning framework that effectively integrates multiview information from both And-Inverter Graphs (AIGs) and Post-Mapping (PM) netlists. At its core, DeepCell employs a self-supervised Mask Circuit Modeling (MCM) strategy, inspired by masked language modeling, to fuse complementary circuit representations from different design stages into unified and rich embeddings. To our knowledge, DeepCell is the first framework explicitly designed for PM netlist representation learning, setting new benchmarks in both predictive accuracy and reconstruction quality. We demonstrate the practical efficacy of DeepCell by applying it to critical EDA tasks such as functional Engineering Change Orders (ECO) and technology mapping. Extensive experimental results show that DeepCell significantly surpasses state-of-the-art open-source EDA tools in efficiency and performance.
00footnotetext: Both authors contributed equally to this research.
@@ I Introduction
Developing Integrated Circuits (ICs) is a highly intricate process that implements complex Boolean logic using billions of transistors, and then precisely places them on a tiny silicon wafer. To manage this complexity, Electronic Design Automation (EDA) tools break down the chip design process into distinct stages, each addressing different representations or modalities of the circuit. These modalities—such as Verilog code, And-Inverter Graphs (AIGs), Post-Mapping (PM) netlists, floorplans, and layouts—capture various aspects of the design at different points in the flow. As such, the EDA process is often viewed as a multimodal transformation [1, 2], where multiple views of the circuit are generated and refined throughout the design cycle.
While emerging circuit learning models [3, 4, 5, 6, 7, 8, 9, 10] have made significant strides in learning representations for individual circuit modalities, they often overlook the multiview information inherent in the design process. As a result, these techniques tend to focus on isolated design tasks at individual stages, such as design-for-test [11], logic optimization [12] and functional reasoning [6, 7], without considering the more diverse spectrum of circuit modalities that could enhance their performance and generalization ability. Although recent cross-modal circuit learning approaches have successfully predicted Power, Performance and Area (PPA) metrics based on the early-stage designs [13, 14], they still face significant challenges in automation and scalability when applied to cross-stage EDA tasks. These tasks, such as logic synthesis, technology mapping and engineering change, requiring integrated insights from multiple modalities remains unexplored. Consequently, how to efficiently leverage the multiview nature of the EDA flow into EDA tools is still an open question.
@@
To answer such critical question, we introduce DeepCell, a self-supervised multiview fusion framework for learning netlist representation. DeepCell integrates information from both pre-mapping (i.e. AIGs) and post-mapping (i.e. PM netlists) modalities, using the corresponding Graph Neural Network (GNN)-based encoders. To fuse these multiview representation, we employ a self-supervised Mask Circuit Modeling (MCM) pretraining task, inspired by Masked Language Modeling (MLM) [15], which leverages embeddings from one view to reconstruct embeddings of another. By bridging the gap between technology-independent and technology-dependent circuit representations using Transformer, DeepCell generates rich, generalizable and multiview embeddings for both AIGs and PM netlists.
We validate DeepCell by applying it to two representative and practical EDA tasks: 1) Functional Engineering Change Orders (ECO): a PM netlist is patched to match a given golden design, avoiding the time-consuming process of re-implementation. 2) Technology mapping: an AIG is converted into PM netlist using a specific technology library. While these tasks may initially seem unrelated, both require a deep understanding of the two circuit views, a challenge that previous circuit learning solutions struggle to address efficiently. DeepCell, with its ability to capture and fuse multiview information, can be seamlessly integrated into existing EDA tools with minimal engineering effort and tiny runtime overhead. In the ECO task, DeepCell significantly reduces both engineering change costs and overall runtime compared to the champion solution [16] of the ICCAD Contest. Besides, DeepCell achieves lower area and delay than the state-of-the-art technology mapping engine in the open-source synthesis tool ABC [17], demonstrating its superior performance and efficiency.
Our key contributions are summarized as follows:
  • We propose DeepCell, a multiview representation learning and fusion framework for circuit netlists, incorporating the information from both technology-independent and technology-dependent circuit designs.
  • We introduce the first representation learning model tailored for PM netlist, employing a self-supervised pretraining mechanism to refine PM netlist embeddings using AIG representations.
  • We demonstrate the effectiveness and generalization ability of DeepCell in various cross-modal EDA tasks, such as functional ECO and technology mapping. DeepCell can be seamlessly integrated as a plug-in into existing EDA tools, achieving significant efficiency improvements over state-of-the-art solutions.
@@
The remainder of this paper is organized as follows: Section II reviews related work. Section III describes the proposed DeepCell framework, including architecture and pretraining mechanism. Section IV presents the pretraining results and investigate the effect of proposed training strategy. Next, we apply DeepCell in functional ECO (see Section V) and technology mapping (see Section VI). Final Section VII concludes this paper.
Figure 1: Overview of DeepCell illustrating the multiview fusion of AIG and PM netlist embeddings
@@ II-A Circuit Representation Learning
Circuit representation learning has emerged as an attractive direction in the field of EDA, focusing on training models to obtain general circuit embeddings that can be applied to various downstream tasks [1]. The first circuit representation learning framework, DeepGate [3], proposes supervising model training using logic-1 probability under random simulation and achieves substantial improvements in tasks like testability analysis [11] and logic equivalence checking [18]. Its successor, DeepGate2 [5], further refines this approach by separating functional and structural embeddings for different applications. Additionally, Gamora [6] and HOGA [7] leverage sub-circuit identification as pre-training tasks, while FGNN [4] trains models in an unsupervised contrastive manner by distinguishing between equivalent and non-equivalent circuits. DeepGate3 [8] and DeepGate4 [19] introduce the framework that combines Graph Neural Networks (GNNs) and Transformer blocks to address the scalability challenges in circuit representation learning, allowing the model to improve with the growth of the training dataset.
While these methods are significant, they primarily focus on single-modal circuit representations (e.g., AIGs), thus limiting their ability to exploit complementary information across multiple circuit modalities. Recent works [13, 14] improve large language models (LLMs) by enhancing the understanding of Verilog code using textual summaries and netlists, but they are not directly compatible with practical EDA tools for circuit-based applications. In contrast, DeepCell framework combines the information from both AIG and PM netlist, fusing circuit representations from different circuit instances. As a result, it stands out in its ability to handle critical EDA tasks that previous approaches struggle to address effectively due to their lack of multiview perception.
@@ II-B Functional ECO
ECO are a critical component in the VLSI design process, used to rectify design problems after tape-out or chip fabrication. ECO involves making modifications to correct these errors, and they are indispensable in avoiding the high expenses associated with design re-spin  [20]. For functional ECO, the purpose is to generate patch so that the given circuit is equal to the golden circuit, while minimizing the resource cost of the generated patches and making the running time as short as possible. Synthesis-based ECO algorithms are good at solving this problem [21]. It relies on a diagnostic strategy to identify internal rectifier signals, and then applies a resynthesis technique to generate patch functions for functional differences. These algorithms have been able to automate the process of functional ECO. However, their effectiveness depends heavily on the quality of the candidate signals for patch generation. When the candidate set is large, the process of selecting rectification cones often results in significant time consumption due to repeated invocations of satisfiability (SAT) and quantified Boolean formula (QBF) solving [22]. Recent approaches attempt to mitigate this issue by prioritizing candidates based on patch cost metrics alone [23, 24]. Yet, these methods overlook driving relations between signals, lead to the incomplete search space exploration. Our DeepCell framework, with its multiview perception, can predict the driven logic of a given cell under the golden implementation. This ability allows DeepCell to reduce the candidate signal set, thereby enhancing the efficiency of ECO tools.
II-C Technology Mapping
Technology mapping is a critical step in logic synthesis, where an optimized Boolean network is transformed into a circuit composed of primitive elements supported by the target technology library. Currently, the most widely used approach for technology mapping is cut-based technology mapping [25]. This method involves enumerating a set of structural cuts for each node in the network, computing the Boolean function of each cut, and matching these functions with the supergates available in the technology library. Matches are sorted based on a target heuristic, such as area, delay, or power consumption, and the optimal cover is selected during a graph traversal [26]. However, this method relies on a load-independent abstraction of delay information, which approximates the delay at the output of a gate using delays at each of its inputs based on a gain-based approach [27]. While this abstraction simplifies the mapping process, it may result in suboptimal mapping quality [28]. Recent works have applied reinforcement learning to technology mapping. For instance, MapTune [29] dynamically optimizes cell library selection in ASIC technology mapping using reinforcement learning, reducing the search space and improving mapping quality. Ye et al. [30] propose a reinforcement learning framework that optimizes timing performance under error constraints to mitigate the effects of aging and variation. These approaches still rely on the fundamental mapping engines with load-independent models. Our DeepCell framework, by fusing information from pre- and post-mapping views, predicts actual loads, leading to more accurate delay predictions and improving the overall mapping quality.
@@ III-A Overview
Fig. 1 presents the overview of DeepCell framework, which consists of a GNN-based AIG encoder, a GNN-based PM encoder and Transformer blocks. The framework operates in two stages to capture multiview information to enhance both AIG and PM netlist representations. In Stage 1, both the AIG and PM encoders are independently trained to extract AIG embeddings (AIG Emb.) and PM netlist embeddings (PM Emb.) from their respective input formats. In Stage 2, the embeddings from both views are concatenated into a token sequence, after which a random subset of tokens is masked. The Transformer blocks then reconstruct the masked tokens, effectively injecting multiview information into the node-level representations.
@@ III-B1 Node Features
PM netlists consist of a wide variety of standard cells, making it impractical to represent them using one-hot encoding for each type of cell. Therefore, to address this, we embed the truth table of each standard cell into its corresponding node feature $x_{i}$. Formally, the node feature encoding is defined in Eq. (1), where $D$ is the dimension of node feature vector and $tt_{i}$ represents the 0/1 truth table vector of standard cell $i$. The truth table $tt_{i}$ is repeated until the node feature vector reaches the specified dimension $D$. In the default setting, we assign $D=64$, ensuring that this encoding mechanism is adaptable to PM netlists across various technology libraries and supports arbitrary logic units with up to 6 inputs, which is typical for most cells.
@@
For example, the cell xor2_1 defines XOR functionality with 2 inputs and 1 output. Its truth table, 0110, is extracted from the technology library and expanded into a 64-dimensional feature vector by repeating the pattern. Thus, the node feature of the xor2_1 cell becomes 0110 0110 ... 0110 $\in\{0,1\}^{64}$. This encoding mechanism can therefore handle arbitrary standard cells by representing their truth tables as binary vectors of a fixed length.
@@ III-B2 Aggregator
For structural embedding aggregation, $aggr^{s}$ is implemented using the GCN aggregator [31], which aggregates messages from the predecessors of $i$. The initial structural embedding of each cell is randomly assigned and considered as its unique identifier. Here, $\mathcal{P}(i)$ denotes the set of fan-in cells of $i$:
@@
For functional embedding aggregation, $aggr^{f}$ is implemented using a self-attention aggregator to distinguish the functionality of the predecessors. We discuss the effectiveness of different aggregators in Sec. IV-B2. Unlike AIG netlists, which consist solely of AND gates and inverters, PM netlists contain diverse standard cells. To account for this diversity, we differentiate cells using their node features $x_{i}$ and introduce an update function, $update$. Formally, the functional aggregation process in DeepCell is defined as Eq. (3), with the $update$ function simply implemented as a multi-layer perceptron (MLP).
@@
Finally, the embeddings of PM netlist are denoted as Eq. (4), where $\Phi^{P}$ is the above GNN-based PM encoder.
$\begin{split}\mathbf{H}^{P}&=\Phi^{P}(\mathcal{G}^{P})\\ H_{i}^{P}&\in\mathbf{H}^{P},i\in\mathcal{V}^{P}\end{split}$ (4)
@@ III-C AIG Encoder
Our DeepCell framework incorporates a multiview representation learning mechanism, allowing it to learn cell embeddings in PM netlists from an additional perspective provided by AIGs and refine AIG embeddings with complementary information from the PM netlists. As shown in Fig. 1, given a PM netlist $\mathcal{G}^{P}=(\mathcal{V}^{P},\mathcal{E}^{P})$, we convert it into an AIG $\mathcal{G}^{A}=(\mathcal{V}^{A},\mathcal{E}^{A})$ using EDA tools. In our implementation, we use the ABC tool [17] with the command ‘strash’ to transform the PM netlist into an AIG through one-level structural hashing. We then utilize DeepGate2 [5], a well-trained and widely-used model for AIG representation learning, as the AIG encoder $\Phi^{A}$ to derive gate-level embeddings $H_{j}^{A}\in\mathbf{H}^{A}$, as defined in Eq. (5). Other AIG encoders, such as PolarGate [10], DeepGate3 [8] and HOGA [7], are also compatible with this framework and are discussed in Section IV-B3. In our default setting, we opt for DeepGate2 [5] as the backbone AIG encoder due to the trade-off between model complexity and performance.
$\begin{split}\mathbf{H}^{A}&=\Phi^{A}(\mathcal{G}^{A})\\ H_{j}^{A}&\in\mathbf{H}^{A},j\in\mathcal{V}^{A}\end{split}$ (5)
@@ III-D Mask Circuit Modeling
Figure 2: Mask circuit modeling of PM netlists
To incorporate multiview information into the GNN-based encoders, we propose the Masked Circuit Modeling (MCM) training mechanism. Unlike the conventional MLM in LLMs [15], where certain tokens are randomly masked and then reconstructed using the remaining tokens, DeepCell masks circuit cones as illustrated in Fig. 2. To be specific, given a PM netlist $\mathcal{G}^{P}$ and the corresponding AIG $\mathcal{G}^{A}$, we randomly select node $p\in\mathcal{G}^{P}$ and extract a $k$-hop predecessors around node $p$, denoted as the Masked Hop $\mathcal{M}(p)$. All nodes $i\in\mathcal{M}(p)$ are considered as masked, and their functional embeddings $hf_{i}$ are replaced by the learnable masked token $hm$, while their structural embeddings $hs_{i}$ are preserving.
As a result, we have $M+N$ tokens (see Eq. (6)) in total, where $M=|\mathcal{V}^{P}|$ and $N=|\mathcal{V}^{A}|$. All the selected masked nodes $p$ is in set $\mathcal{S}$.
$\begin{split}{H_{i}^{P}}&=\{hs_{i}^{P},hf_{i}^{P}\},~{}i\in\mathcal{V}^{P},i\notin\mathcal{S}\\ {H_{i}^{P}}&=\{hs_{i}^{P},hm\},~{}i\in\mathcal{V}^{P},i\in\mathcal{S}\\ {H_{j}^{A}}&=\{hs_{j}^{A},hf_{j}^{A}\},~{}j\in\mathcal{V}^{A}\\ \end{split}$ (6)
Then, we process these $M+N$ tokens using a Transformer model $\mathcal{T}$. Formally, we define the input and output of the Transformer model as below.
@@
Similarly, when the DeepCell framework is applied to refine the AIG embeddings, the MCM mechanism generates masked AIG tokens and non-masked PM tokens. It is important to note that recovering the node type is unnecessary, as it can be directly inferred by counting the fan-in degree in the AIG, where an AND gate always has 2 fan-in pins and a NOT gate has 1 fan-in pin. Consequently, the training objective is to reconstruct the functional embeddings using information from surrounding nodes and nodes in the other modality.